# Bourbaki degree of pairs of projective surfaces

### 31 Mar 2026 v3

@@ 8 $e\leq 3$
so that $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{3}}=3$.
For the prime $\mathfrak{p}_{4}=(x_{0},x_{2})$, we take $u=(x_{3}(2x_{2}-x_{1}))^{-1}\in R_{\mathfrak{p}_{4}}$ and the elementary operations $R_{1}^{\prime}=R_{1}-u[-x_{0}^{2}+2x_{2}(x_{1}-x_{0}+x_{3})]R_{2}\doteq R_{1}-uQR_{2}$ and $R_{2}^{\prime}=uR_{2}$, to obtain:
@@
where the elements $(x_{2}-x_{1}),(x_{1}-x_{0}+x_{3}),x_{3}\in R_{\mathfrak{p}_{4}}^{\times}$. Since $f_{1}\in I$, we have
$x_{2}^{2}\equiv 2x_{0}(x_{1}-x_{2})\mod I,$
and thus
@@
with $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{4}}=2$. Using the associativity formula to compute $m(\sigma)$, we obtain:
@@
On the other hand, $e(\sigma)=2$, so using the Bourbaki formula we obtain $\operatorname{Bour}(\sigma)=1$, and conclude also the minimal free resolution for ${\mathcal{T}}_{\sigma}$ of the form
@@
  • $e=2$ and $\sigma$ is a nearly free sequence with $c_{3}({\mathcal{T}}_{\sigma})=2$ (see 41);
@@
From 32, we know that $e\leq 3$. For $\mu$-semistable sheaves, $e\geq 2$ and, since $c_{2}({\mathcal{T}}_{\sigma}(2))=1$ and $c_{1}({\mathcal{T}}_{\sigma}(2))=0$, from [undefb, Lemma 2.1], we conclude that the only possible cases are $c_{3}=2$ (strictly semistable case) or $c_{3}=0$ (stable case). For the strictly semistable case $e=2$ we obtain $\operatorname{Bour}(\sigma)=1$, so we get the first case.
For the second case, we must have a stable bundle ${\mathcal{T}}_{\sigma}(2)$ with Chern classes $(0,1,0)$, which are precisely null correlation bundles described in [undefx] fitting in a sequence of the form
@@
    There is only one case of nearly free sequence $\sigma$, with discrete invariants $m(\sigma)=7$, $e=2$ and $c_{3}({\mathcal{T}}_{\sigma})=2$ (see Example 42), which is strictly $\mu$-semistable;
@@
Item $(a)$ is 34 and item $(b)$ with 37 and Example 40. Item $(c)$ follows from 43 and item $(d)$ is 36. ∎
To finish this study, we consider a strictly $\mu$-semistable pencil of cubics with $m(\sigma)=4$.
@@
Considering the following pencil of cubics $(d_{f}=d_{g}=2)$, where the first one is smooth:
@@
The scheme structures $\operatorname{supp}(\mathcal{Q}_{\sigma})=\Xi_{\sigma}$ coincide, with support at a line $V(x_{2},x_{3})$. Over the prime ideal $\mathfrak{p}=(x_{2},x_{3})$, we perform some elementary operations for the matrix $\nabla\sigma$, first $R_{1}^{\prime}=R_{1}-R_{2}$, and then $R_{2}^{\prime}=u/3\cdot R_{2}$, where $u=(x_{1}^{2})^{-1}\in R_{\mathfrak{p}}$, to obtain the form:
@@
$(\mathcal{Q}_{\sigma})_{\mathfrak{p}}$ $\simeq\frac{R_{P}}{(3x_{2}^{2}-x_{3}^{2},3x_{3}^{2}-2x_{2}x_{3})}$
@@
and thus we conclude $\operatorname{{length}}(\mathcal{Q})_{\mathfrak{p}}=4$. Since $\deg(R/\mathfrak{p})=1$, using the associativity formula we obtain $m(\sigma)=4$. A first syzygy is $\nu=(-x_{1}^{2},x_{0}^{2},0,0)^{T}$, so we obtain $e=2$, and thus $\operatorname{Bour}(\sigma)=4$.
@@
To finish this subsection, we explore in an example that the Bourbaki degree of a pencil of cubics does not only depend on the singularity type of each cubic, but rather how these align when we consider the associated pencil.
@@ 4.2 Degree 6 curves inside quadric surfaces
We start by applying Theorem 28 to the sheaf $\mathcal{E}={\mathcal{T}}_{\sigma}$, assuming $\sigma$ is incompressible. The quantities of the theorem become
$A$ $=\left\lceil\frac{1}{3}(c_{1}^{2}+2c_{1}+3)\right\rceil=\left\lceil\frac{1}{3}(9-6+3)\right\rceil=0$
$B$ $=\left\lceil\frac{1}{3}(c_{1}^{2}+3c_{1}+8)\right\rceil-1=\left\lceil\frac{1}{3}(9-9+8)\right\rceil-1=1.$
Moreover, another important invariant is: $c_{2}(\mathcal{E})=7-m(\sigma)$. Direct application of the result shows the next pair of propositions:
@@
If $\sigma=(f,g)$ is an incompressible normal sequence with $d_{f}=1,d_{g}=2$, then $m(\sigma)\leq 5$.
Proof.
Let us assume that $e\in[1,d-2]$ and $d\geq 3$. Then, from the Bourbaki formula and the fact that $\operatorname{Bour}(\sigma)\geq 0$, we obtain
$m(\sigma)\leq m_{0}+e(e-d)\leq m_{0}+e-d\leq m_{0}-2,$
which gives the claim. This covers the case $e=1$. If we now assume $e>1$ and apply Theorem 28 to $\mathcal{E}={\mathcal{T}}_{\sigma}(1)$, a reflexive sheaf with $c_{1}=-1$, we obtain the constant is
$A$ $=\left\lceil\frac{1}{3}(c_{1}^{2}+2c_{1}+3)\right\rceil=\left\lceil\frac{1}{3}(1-2+3)\right\rceil=0.$
Hence, $5-m(\sigma)=c_{2}({\mathcal{T}}_{\sigma}(1))=c_{2}(\mathcal{E})\geq A=0$ and thus $m(\sigma)\leq 5$. ∎
Proposition 47.
Let $\sigma=(f,g)$ be an incompressible normal sequence with $d_{f}=1$ and $d_{g}=2$. Then, the following hold:
@@
    If $6\leq m(\sigma)\leq 7$, then $c_{3}\leq 5-m(\sigma)$.
@@
    If $0\leq m(\sigma)<6$, then $c_{3}\leq m(\sigma)^{2}-14m(\sigma)+47$.
@@
Next, we apply Theorem 29 to the sheaf $\mathcal{E}={\mathcal{T}}_{\sigma}(1)$, a sheaf with $c_{1}=-1$ and $c_{2}(\mathcal{E})=5-m(\sigma)\geq 0$. Then, the theorem says that, for
@@
we have $h^{0}(\mathcal{E}(t))=h^{0}({\mathcal{T}}_{\sigma}(t+1))\neq 0$.
Direct application of this bound for high values of $m(\sigma)=4,5$ we get $t=0,t=1$, respectively. This shows the following proposition.
@@
Let $\sigma=(f,g)$ be a normal sequence with $d_{f}=1$, $d_{g}=2$. Then $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})\leq 3$, and we have the following table of bounds for the following possibilities of $m(\sigma)$:
$m(\sigma)$ $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$
4 $e\leq 2$
5 $e\leq 1$
As an easy consequence of Remark 6, we obtain:
@@ 5 $e\leq 1$
Let $\sigma=(f,g)$ be a normal sequence which is not a pencil, (i.e., with $d_{f}\neq d_{g}$) such that $d$ is odd. If $\operatorname{Bour}(\sigma)$ is odd, then $\sigma$ is not locally free.
@@
Assuming ${\mathcal{T}}_{\sigma}$ is locally free, we get $c_{3}({\mathcal{T}}_{\sigma})=0$, and therefore
$2p_{a}(B)$ $=2-\operatorname{Bour}(\sigma)(4+d-2e)$
$p_{a}(B)$ $=1-\operatorname{Bour}(\sigma)(2+\frac{d}{2}-e)=1-(2-e)\operatorname{Bour}(\sigma)+\frac{d\operatorname{Bour}(\sigma)}{2},$
which is not an integer if $\operatorname{Bour}(\sigma)$ is odd, so we get a contradiction. ∎
@@
To show this, we show that in any of these cases the Bourbaki degree $\operatorname{Bour}(\sigma)$ is odd for every possibility of $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$, and then the result follows from 51.
From 3, we get $\operatorname{Bour}(\sigma)=e^{2}-3e+7-m(\sigma)$, so that when $m(\sigma)$ is even, $7-m(\sigma)$ is odd. We claim $e^{2}-3e$ is always an even number for $e\geq 0$ integer.
Assuming $e=2k$ is even, we obtain
@@
Assuming $\sigma$ is non-compressible, by 50, we obtain that $e=1$. But from the formula for the Bourbaki degree with $e=1$, $m(\sigma)=5$ we obtain $\operatorname{Bour}(\sigma)=0$, and thus $\sigma$ must be free. On the other hand, if we assume ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)$, then $e=1$ and $\operatorname{Bour}(\sigma)=0$ give $m(\sigma)=5$. ∎
@@
Furthermore, these are the only two possibilities of numerical invariants for nearly free sequences with $d_{f}=1,d_{g}=2$.
@@
Using 50, $e\in\{1,2\}$, and the two cases imply $\operatorname{Bour}(\sigma)=1$. Using Remark 6 for $p_{a}(B)=0,\deg(B)=1$, we obtain the $c_{3}$’s above, and both appear as examples. To conclude the last claim, we note that if $m(\sigma)\leq 3$, then
@@
for $e\in\{1,2,3\}$. ∎
Example 56 (Nearly free sequence with $d_{f}=1,d_{g}=2$ and $e=2$).
@@
The irreducible components of $\Xi_{\sigma}$ consist of three lines $V(x_{1},x_{3})$, $V(x_{1},x_{0}-x_{2})$ and $V(x_{3},x_{0}-x_{2})$ and a point $p=V(x_{2},x_{1}-x_{3},x_{0})$ outside the three lines, so $p\in\operatorname{Sing}({\mathcal{T}}_{\sigma})$. We may compute $m(\sigma)=4$, where the first line above occurs with length two and the other two occur with length one.
For $\mathfrak{p}_{1}=(x_{1},x_{3})$, we may consider $u=(x_{0})^{-1}$ and perform two elementary operations, namely $R_{2}^{\prime}=R_{2}-ux_{3}(x_{0}-x_{2})R_{1}$ and $R_{1}^{\prime}=uR_{1}$ to rewrite the matrix:
@@
simplifying the invertible elements $(1-u(x_{0}-x_{2})),(x_{0}-x_{2})\notin\mathfrak{p}_{1}$, hence $\operatorname{{length}}((\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}})=2$.
For $\mathfrak{p}_{2}=(x_{1},x_{0}-x_{2})$, consider $u=(-x_{3})^{-1}\in R_{\mathfrak{p}_{2}}$ and perform two elementary operations $R_{2}^{\prime}=R_{2}+ux_{1}x_{3}R_{1}$ and $R_{1}^{\prime}=uR_{1}$ to rewrite $\nabla\sigma$ as
@@
and thus $I=(x_{0}-x_{2},x_{1})=\mathfrak{p}_{2}R_{\mathfrak{p}_{2}}$, so that $(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}\simeq\kappa$, and thus $\operatorname{{length}}((\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}})=1$.
For $\mathfrak{p}_{3}=(x_{3},x_{0}-x_{2})$, take $u=(x_{1})^{-1}\in R_{\mathfrak{p}_{3}}$ and the elementary operations $R_{2}^{\prime}=R_{2}-ux_{1}x_{3}R_{1}$, $R_{1}^{\prime}=uR_{1}$ to obtain
@@
and thus $I=(x_{3},x_{0}-x_{2})=\mathfrak{p}_{2}R_{\mathfrak{p}_{2}}$, so that $(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}\simeq\kappa$, and thus $\operatorname{{length}}((\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}})=1$. From the associativity formula over $\mathfrak{p}_{1},\mathfrak{p}_{2},\mathfrak{p}_{3}$, we then obtain
$m(\sigma)=2\cdot 1+1\cdot 1+1\cdot 1=4.$
From the Bourbaki degree formula, once we know that $e=2$, we conclude that $\operatorname{Bour}(\sigma)=1$ and this is a nearly free sequence. The complete minimal free resolution of ${\mathcal{T}}_{\sigma}$ can be obtained computationally
@@
which corresponds to the nearly free resolution given in 16. We have $c_{3}({\mathcal{T}}_{\sigma})=1$, corresponding to the point $p$, the unique irreducible component of $\Xi_{\sigma}$ of codimension three.
@@
We point out that ${\mathcal{T}}_{\sigma}(1)$ will be a stable rank two reflexive sheaf of Chern classes $(-1,2,c_{3})$, since $c_{2}({\mathcal{T}}_{\sigma}(1))=5-m(\sigma)$. Then, the three possibilities $c_{3}\in\{0,2,4\}$ imply that $H^{0}({\mathcal{T}}_{\sigma}(2))\neq 0$, and therefore $e=2$. This follows from [undefw, Proposition 1.1] for $c_{3}=0$, [undefb, Lemma 2.4] for $c_{3}=2$ and [undefn, Lemma 9.6] for $c_{3}=4$. ∎
The unstable case of the above Proposition is numerically described in 27.
@@
Over the prime $\mathfrak{p}_{1}=(x_{0},x_{3})$, let us take $u\doteq(x_{0}-x_{1})^{-1}\in R_{\mathfrak{p}_{1}}$ and $Q\doteq x_{0}x_{1}+3x_{3}^{2}$. Consider the elementary operations $R_{2}=R_{2}-u(x_{0}x_{1}+3x_{3}^{2})R_{1}$ and $R_{1}^{\prime}=uR_{1}$, to rewrite
@@
Therefore, since $f_{2}\in I$ and $u(x_{0}x_{1}+3x_{3}^{2})\in\mathfrak{p}_{1}$, we conclude $x_{3}\in I$. Rewriting $uf_{1}$ and removing terms with $x_{3}$, we obtain
$uf_{1}=2x_{0}^{3}x_{1}x_{2},$
and since $f_{3}=x_{0}^{2}\in I$, we obtain $I=(x_{0}^{2},x_{3})$, hence
@@
so that $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}}=2$.
Over $\mathfrak{p}_{2}=(x_{3},x_{0}-x_{1})$, take $u\doteq(x_{0}^{2})^{-1}$ and consider the elementary operation and $R_{2}^{\prime}=uR_{2}$, to obtain the matrix
@@
and therefore $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}=1$.
Thus, from the additivity formula, we obtain $m(\sigma)=3$. Moreover, $e=2$, so that $\operatorname{Bour}(\sigma)=2$.
The sheaf ${\mathcal{T}}_{\sigma}$ admits a free resolution of the form:
@@
Over $\mathfrak{p}_{1}$, we may choose $u=(2x_{1})^{-1}\in R_{\mathfrak{p}_{1}}$ to perform the elementary operations $R_{2}^{\prime}=R_{2}-ux_{3}(x_{3}-x_{2})R_{1}$ and $R_{1}^{\prime}=uR_{1}$, to rewrite:
@@
since $x_{0}-x_{1}$ is invertible. Thus, $I\simeq\mathfrak{p}_{1}R_{\mathfrak{p}_{1}}$ and $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}}=1$.
@@
Since $(x_{2}-2x_{3})$ is invertible in $R_{\mathfrak{p}_{2}}$, we obtain $(x_{0}-x_{1})\in I$. From this fact, since $x_{2}-x_{3}$ is invertible and $f_{1}\in R_{\mathfrak{p}_{2}}$, we obtain $x_{3}\in I$, and thus $I\simeq\mathfrak{p}_{2}R_{\mathfrak{p}_{2}}$ and $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}=1$.
@@
so that $I\simeq\mathfrak{p}_{3}R_{\mathfrak{p}_{3}}$ and $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{3}}=1$. Thus, from the associativity formula $m(\sigma)=3$. Moreover, $e=2$ and thus $\operatorname{Bour}(\sigma)=2$.
Here, the sheaf ${\mathcal{T}}_{\sigma}$ admits a free resolution below
@@
so that $c_{3}({\mathcal{T}}_{\sigma})=4$, and the associated Bourbaki scheme is a plane smooth quadric.

### 27 Apr 2026 v4

@@ 4.1 Pencils of cubics
so that $\operatorname{\operatorname{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{3}}=3$.
For the prime $\mathfrak{p}_{4}=(x_{0},x_{2})$, we take $u=(x_{3}(2x_{2}-x_{1}))^{-1}\in R_{\mathfrak{p}_{4}}$ and perform elementary matrix operations to obtain:
@@
where the elements $(x_{2}-x_{1}),(x_{1}-x_{0}+x_{3}),x_{3}\in R_{\mathfrak{p}_{4}}^{\times}$. Since $f_{1}\in I$, we have $x_{2}^{2}\equiv 2x_{0}(x_{1}-x_{2})\mod I$, and thus
@@
with $\operatorname{\operatorname{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{4}}=2$. Using the associativity formula to compute $m(\sigma)$, we obtain:
@@
On the other hand, $e=2$, so using the Bourbaki formula we obtain $\operatorname{Bour}(\sigma)=1$, and conclude also the minimal free resolution for ${\mathcal{T}}_{\sigma}$ of the form
@@
  • $e=2$ and $\sigma$ is a nearly free sequence with $c_{3}({\mathcal{T}}_{\sigma})=2$ (see 42);
@@
From 33, we know that $e\leq 3$. For $\mu$-semistable sheaves, $e\geq 2$ and, since $c_{2}({\mathcal{T}}_{\sigma}(2))=1$ and $c_{1}({\mathcal{T}}_{\sigma}(2))=0$, from [4, Lemma 2.1], we conclude that the only possible cases are $c_{3}=2$ (strictly semistable case) or $c_{3}=0$ (stable case). For the strictly semistable case $e=2$ we obtain $\operatorname{Bour}(\sigma)=1$, so we get the first case.
For the second case, we must have a stable bundle ${\mathcal{T}}_{\sigma}(2)$ with Chern classes $(0,1,0)$, which are precisely null correlation bundles described in [23] fitting in a sequence of the form
@@
Summarizing the main results for normal pencils of cubics, we get the following Theorem:
@@
    There is only one case of nearly-free sequence $\sigma$, with discrete invariants $m(\sigma)=7$, $e=2$ and $c_{3}({\mathcal{T}}_{\sigma})=2$ (see Example 43), which is strictly $\mu$-semistable;
@@
Item $(a)$ is 35 and item $(b)$ with 38 and Example 41. Item $(c)$ follows from 44 and item $(d)$ is 37. ∎
Next, we consider some examples. First, a strictly $\mu$-semistable pencil of cubics with $m(\sigma)=4$.
@@
Consider the following pencil of cubics $(d_{f}=d_{g}=2)$, where the first one is smooth:
@@
The scheme structures $\operatorname{supp}(\mathcal{Q}_{\sigma})=\Xi_{\sigma}$ coincide, with support at a line $V(x_{2},x_{3})$. Over the prime ideal $\mathfrak{p}=(x_{2},x_{3})$, we perform some elementary operations for the matrix $\nabla\sigma$ to obtain the form:
@@
$(\mathcal{Q}_{\sigma})_{\mathfrak{p}}$ $\simeq\frac{R_{\mathfrak{p}}}{(3x_{2}^{2}-x_{3}^{2},3x_{3}^{2}-2x_{2}x_{3})}$
@@
and thus we conclude $\operatorname{\operatorname{length}}(\mathcal{Q})_{\mathfrak{p}}=4$. Since $\deg(R/\mathfrak{p})=1$, using the associativity formula we obtain $m(\sigma)=4$. A first syzygy is $\nu=(-x_{1}^{2},x_{0}^{2},0,0)^{T}$, so we obtain $e=2$, and thus $\operatorname{Bour}(\sigma)=4$.
@@
To finish this subsection, we explore in an example that the Bourbaki degree of a pencil of cubics does not depend only on the singularity type of each cubic, but rather how these align when we consider the associated pencil.
@@ 4.2 Degree 6 curves inside quadric surfaces
We start by applying Theorem 30 to the sheaf $\mathcal{E}={\mathcal{T}}_{\sigma}$, assuming $\sigma$ is incompressible. The quantities become $A=B=2$. Moreover, $c_{2}(\mathcal{E})=7-m(\sigma)$, so direct application of the result shows the following:
@@
Let $\sigma=(f,g)$ be an incompressible normal sequence with $d_{f}=1$ and $d_{g}=2$. Then $m(\sigma)\leq 5$ and the following hold:
@@
    If $m(\sigma)=5$, then $c_{3}({\mathcal{T}}_{\sigma})=0$.
@@
    If $0\leq m(\sigma)<4$, then $c_{3}\leq m(\sigma)^{2}-12m(\sigma)+35$.
@@
Next, we apply Theorem 31 to the sheaf $\mathcal{E}={\mathcal{T}}_{\sigma}(1)$, a sheaf with $c_{1}=-1$ and $c_{2}(\mathcal{E})=5-m(\sigma)\geq 0$. Then, for
@@
we have $h^{0}(\mathcal{E}(t))=h^{0}({\mathcal{T}}_{\sigma}(t+1))\neq 0$. Direct application of this bound shows the following proposition.
@@
Let $\sigma=(f,g)$ be a normal sequence with $d_{f}=1$, $d_{g}=2$. Then $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})\leq 3$. Whenever $m(\sigma)=4$, $e\leq 2$ and, whenever $m(\sigma)=5$, $e\leq 1$.
We explore some parity results in the next two propositions about local freeness.
@@
Let $\sigma=(f,g)$ be a normal sequence such that $d=d_{f}+d_{g}$ is odd. If $\operatorname{Bour}(\sigma)$ is odd, then $\sigma$ is not locally free.
@@
Assuming ${\mathcal{T}}_{\sigma}$ is locally free, we get $c_{3}({\mathcal{T}}_{\sigma})=0$, and therefore $2p_{a}(B)-2=-\operatorname{Bour}(\sigma)(4+d-2e)$. But the left-hand-side is even, and the right-hand side is a product of two odd numbers, hence odd, so we get a contradiction. ∎
@@
We show that, in any of these cases above, the Bourbaki degree $\operatorname{Bour}(\sigma)=e^{2}-3e+7-m(\sigma)$ is odd, thus the result follows from 51. Whenever $m(\sigma)$ is even, $7-m(\sigma)$ is odd. We claim $e^{2}-3e$ is always an even number for $e\geq 0$, and therefore $\operatorname{Bour}(\sigma)$ is always odd for $m(\sigma)$ even. First, assuming $e=2k$ is even, we obtain
@@
Assuming $\sigma$ is incompressible, by 50, we obtain that $e=1$. But from the formula for the Bourbaki degree with $e=1$, $m(\sigma)=5$ we obtain $\operatorname{Bour}(\sigma)=0$, and thus $\sigma$ must be free. On the other hand, if we assume ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)$, then $e=1$ and $\operatorname{Bour}(\sigma)=0$ give $m(\sigma)=5$. ∎
@@
Furthermore, these are the only two possibilities of numerical invariants for nearly-free sequences with $d_{f}=1,d_{g}=2$.
@@
Using 50, $e\in\{1,2\}$, and both cases imply $\operatorname{Bour}(\sigma)=1$ when $m(\sigma)=4$. Using $p_{a}(B)=0$ and $\deg(B)=1$, we obtain the $c_{3}$’s above, and both can be realized. To conclude the last claim, for $m(\sigma)\leq 3$,
@@
for possible values $1\leq e\leq 3$. ∎
Example 56 (Nearly-free sequence with $d_{f}=1,d_{g}=2$ and $e=2$).
@@
The irreducible components of $\Xi_{\sigma}$ consist of three lines $V(x_{1},x_{3})$, $V(x_{1},x_{0}-x_{2})$ and $V(x_{3},x_{0}-x_{2})$ and a point $p=V(x_{2},x_{1}-x_{3},x_{0})$ outside the three lines. We compute $m(\sigma)=4$, where the first line above occurs with multiplicity two and the other two are simple.
For $\mathfrak{p}_{1}=(x_{1},x_{3})$, we may consider $u=(x_{0})^{-1}$ and perform elementary operations to rewrite the matrix as
@@
simplifying the invertible elements $(1-u(x_{0}-x_{2})),(x_{0}-x_{2})\notin\mathfrak{p}_{1}$, hence $\operatorname{\operatorname{length}}((\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}})=2$.
For $\mathfrak{p}_{2}=(x_{1},x_{0}-x_{2})$, consider $u=(-x_{3})^{-1}\in R_{\mathfrak{p}_{2}}$ and perform elementary operations to rewrite $\nabla\sigma$ as
@@
and thus $I=(x_{0}-x_{2},x_{1})=\mathfrak{p}_{2}R_{\mathfrak{p}_{2}}$, so that $(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}\simeq\kappa$, and thus $\operatorname{\operatorname{length}}((\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}})=1$.
For $\mathfrak{p}_{3}=(x_{3},x_{0}-x_{2})$, take $u=(x_{1})^{-1}\in R_{\mathfrak{p}_{3}}$ and an elementary operations to obtain
@@
and thus $I=(x_{3},x_{0}-x_{2})=\mathfrak{p}_{2}R_{\mathfrak{p}_{2}}$, so that $(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}\simeq\kappa$, and thus $\operatorname{\operatorname{length}}((\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}})=1$. From the associativity formula over $\mathfrak{p}_{1},\mathfrak{p}_{2},\mathfrak{p}_{3}$, we then obtain $m(\sigma)=2\cdot 1+1\cdot 1+1\cdot 1=4$.
From the Bourbaki degree formula, once we know that $e=2$, we conclude that $\operatorname{Bour}(\sigma)=1$ and this is a nearly-free sequence. The minimal free resolution of ${\mathcal{T}}_{\sigma}$ can be obtained computationally
@@
which corresponds to the nearly-free resolution given in 17. We have $c_{3}({\mathcal{T}}_{\sigma})=1$, corresponding to the point $p$, the unique irreducible component of $\Xi_{\sigma}$ of codimension three.
@@
The sheaf ${\mathcal{T}}_{\sigma}(1)$ is a stable rank two reflexive sheaf with Chern classes $(-1,2,c_{3})$, and $c_{3}\leq c_{2}^{2}=4$, using [14, Theorem 8.2]. All three possibilities $c_{3}\in\{0,2,4\}$ imply $e=2$. Indeed, this follows from [22, Proposition 1.1] for $c_{3}=0$, [4, Lemma 2.4] for $c_{3}=2$ and [14, Lemma 9.6] for $c_{3}=4$. ∎
Next examples illustrate some cases of the proposition above.
@@
Over the prime $\mathfrak{p}_{1}=(x_{0},x_{3})$, let us take $u\doteq(x_{0}-x_{1})^{-1}\in R_{\mathfrak{p}_{1}}$ and $Q\doteq x_{0}x_{1}+3x_{3}^{2}$. Consider elementary operations to rewrite
@@
Therefore, since $f_{2}\in I$ and $u(x_{0}x_{1}+3x_{3}^{2})\in\mathfrak{p}_{1}$, we conclude $x_{3}\in I$. Rewriting $uf_{1}$ and removing terms with $x_{3}$, we obtain $uf_{1}=2x_{0}^{3}x_{1}x_{2}$, and since $f_{3}=x_{0}^{2}\in I$, we obtain $I=(x_{0}^{2},x_{3})$, hence
@@
so that $\operatorname{\operatorname{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}}=2$.
Over $\mathfrak{p}_{2}=(x_{3},x_{0}-x_{1})$, take $u\doteq(x_{0}^{2})^{-1}$ and consider an elementary operation to obtain a matrix
@@
and therefore $\operatorname{\operatorname{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}=1$. Thus, from the additivity formula, we obtain $m(\sigma)=3$. Moreover, $e=2$, so that $\operatorname{Bour}(\sigma)=2$. The sheaf ${\mathcal{T}}_{\sigma}$ admits a free resolution of the form:
@@
Over $\mathfrak{p}_{1}$, we may choose $u=(2x_{1})^{-1}\in R_{\mathfrak{p}_{1}}$ to perform the elementary operations and rewrite:
@@
since $x_{0}-x_{1}$ is invertible. Thus, $I\simeq\mathfrak{p}_{1}R_{\mathfrak{p}_{1}}$ and $\operatorname{\operatorname{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}}=1$.
@@
Since $(x_{2}-2x_{3})$ is invertible in $R_{\mathfrak{p}_{2}}$, we obtain $(x_{0}-x_{1})\in I$. From this fact, since $x_{2}-x_{3}$ is invertible and $f_{1}\in R_{\mathfrak{p}_{2}}$, we obtain $x_{3}\in I$, and thus $I\simeq\mathfrak{p}_{2}R_{\mathfrak{p}_{2}}$ and $\operatorname{\operatorname{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}=1$.
@@
so that $I\simeq\mathfrak{p}_{3}R_{\mathfrak{p}_{3}}$ and $\operatorname{\operatorname{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{3}}=1$. Thus, from the associativity formula $m(\sigma)=3$. Moreover, $e=2$ and thus $\operatorname{Bour}(\sigma)=2$. Here, the sheaf ${\mathcal{T}}_{\sigma}$ admits a free resolution below
@@
so that $c_{3}({\mathcal{T}}_{\sigma})=4$, and the associated Bourbaki scheme is a plane conic.